OpenStudy (anonymous):

Shalika bought a purse for $120. The tax rate is 9%. Which proportion can be used to find the amount of tax Shalika will pay? a.9/120=x/100 b.120/x=100/9 c.x/120=100/9 d.120/9=100/x

OpenStudy (anonymous):

http://www.mathgoodies.com/lessons/percent/proportions.html see if you can get it from here. if not, i can help you from there

OpenStudy (anonymous):

@lordoftanks do you understand?

OpenStudy (anonymous):

i guess help would be better :)

OpenStudy (anonymous):

yes i do but this one is harder than the questions i do

OpenStudy (anonymous):

It's x/120 = 100/9. You would then have to cross multiply (100x=1080) and divide each by 100 to get x=10.8, do you understand why? x is how much tax she would have to pay

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so x/12=100/9 is the answer?

OpenStudy (anonymous):

yes